A car is capable of accelerating at a rate of 3.20 m/s^2. If it is going 11.0 m/s when it leaves the curved portion of the entrance ramp to an expressway, how much straightaway does it need to obtain 35.0 m/s before merging with the traffic ?
Answer by fractalier(6550)

We can use the equation
Vf^2 - Vo^2 = 2ax
to find the distance needed, x.
x = (35^2 - 11^2) / 2(3.2)
x = (1225 - 121) / 6.4
x = 172.5 m